If 316 mL nitrogen is combined with 178 mL oxygen, what volume of N2O is produced at constant temperature and pressure if the reaction proceeds to 82.0% yield? 2N2(g) + O2(g) → 2N20(g) mL
Answer #1

Balanced chemical equation is:

2 N2 + O2 ---> 2 N2O

2 mL of N2 reacts with 1 mL of O2

for 3.16*10^2 mL of N2, 1.58*10^2 mL of O2 is required

But we have 1.78*10^2 mL of O2

so, N2 is limiting reagent

we will use N2 in further calculation

According to balanced equation

volume of N2O formed = (2/2)* volume of N2

= (2/2)*3.16*10^2

= 3.16*10^2 mL

This is theoretical yield

Now use:

% yield = actual yield * 100 / theoretical yield

82.0 = actual yield * 100 / 3.16*10^2

actual yield = 259 mL

Answer: 259 mL